OpenStudy (anonymous):

What is 20% of 12,000? 240 2,400 1,200 120 @melanielovee4

OpenStudy (anonymous):

2400

OpenStudy (anonymous):

Tell me how you get the answer woman!

OpenStudy (anonymous):

You answer so fast I don't even know if it's correct or not.

OpenStudy (anonymous):

oh okay lolol

OpenStudy (anonymous):

I'm sorry I'm just trying not to get the wrong answer right now cause this is like a big test for this online class I take and it's such a pain...

OpenStudy (anonymous):

so if you want to get something percent, just divide by 100 and multiply however many percent

OpenStudy (anonymous):

its okay good luck!

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

For example, how to solve for a value when you know what a percentage part is: 15 is 10% what number? in equation form is 15/10% = X. We need to first convert the percentage to an equivalent decimal to do our calculation. Converting a percentage to a decimal we remove the percentage sign and divide by 100.

OpenStudy (mathstudent55):

In Math, "of" usually means "times". Also, "is" means "=". Also, "%" means "/100" (or "divide by 100"). Now translate your problem from English to Math: What is 20% of 12,000? ? = 20% * 12,000 ? = 20/100 * 12,000 ? = 0.2 * 12,000 Now multiply 0.2 by 12,000 to find the answer.